OpenStudy (joshoyen):

Find all roots. 1. x^4 - 16x^2 + 63 = 0

OpenStudy (snowsurf):

I won't solve it for you. But lets pick a variable to substitute. Let \[s = x ^{2}\] and you will have the following equation \[s ^{2}-16s+63=0\] Factor as you normally would then once you done then substitute back x^2 term.

OpenStudy (joshoyen):

s=7,s=9

OpenStudy (snowsurf):

Very good. So you should have (s-9)(s-7) = 0. Now put back in the x^2 term.

OpenStudy (joshoyen):

(s^2-9)(s^2-7)?

OpenStudy (snowsurf):

Yes and set it to zero. Can you factor any more? Take a look at it.

OpenStudy (joshoyen):

hm im not sure on that one

OpenStudy (whpalmer4):

@joshoyen you should always think "possible difference of squares" when you see something like \[x^2 - \text{something}\] remember that the formula for a difference of squares is \[(x-a)(x+a) = x^2 +ax -ax - a^2 = x^2 -a ^2\]

OpenStudy (snowsurf):

whpalmer4 has a good template for you to figure that last part. Look at the (s^2-9) term.